OpenStudy (anonymous):

how many numbers from 0 to 999 are not divisible by either 5 or 7............?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

20 are divisible by 5 14 are divisible by 7 leaving 100 - 20 -14 = 66 66 numbers not divisible by 5 or 7

OpenStudy (anonymous):

dumbcow, there are 1000 numbers from 0-999

OpenStudy (dumbcow):

haha oops didnt see enough 9's

OpenStudy (dumbcow):

then its 200 for 5 142 for 7 leaving 1000 -200-142 = 658 658 numbers

OpenStudy (anonymous):

@dumbcow: you have to consider the numbers that are divisible by both \(5\) and \(7\) and subtract them from the total.

OpenStudy (dumbcow):

wow, im really off ok like 35,70... there are 28 numbers divisible by 35 658 - 28 = 630

OpenStudy (anonymous):

i think there are 143 divisible by 7. 7*142=994 and then 0 also

OpenStudy (dumbcow):

i meant 658 +28 = 686

OpenStudy (anonymous):

#s divisible by 5=200 #s divisible by 7=143 #divisible by 7 and 5=29 So the number of numbers in [0,999] that are not divisible by 5 or 7 is 1000-200-143+29=686